OpenStudy (anonymous):

what is 5√7+3√17+6√17-8√7+2√7, with the work please.

OpenStudy (anonymous):

\[5\sqrt{7}-8\sqrt{7}+2\sqrt{7}+3\sqrt{17}+6\sqrt{17}\] put like terms together

OpenStudy (nincompoop):

this question is too easy. it's beneath you

OpenStudy (anonymous):

\[-\sqrt{7}+9\sqrt{17}\]